COMMISSION 6 : ASTRONOMICAL TELEGRAMS (TELEGRAMMES ASTRONOMIQUES)

Report of Meeting, 23 August 1997

President: R. M. West Secretary: B. G. Marsden

After the adoption of the agenda for the meeting, President West welcomed the eight other persons present and asked them to stand in memory of three members lost since the last meeting: Michael P. Candy, Antonín Mrkos and Leonida Rosino. Remarking that Commission 6 is one of the IAU’s earliest and that the rapid dissemination of urgent astronomical information by the Central Bureau for Astronomical Telegrams (CBAT) is one of the oldest and most essential services of the IAU, he reminded those present that one of his hopes as president had been that it would be possible to disseminate the IAU Circulars free of charge: after all, in some countries, the cost of a subscription was a sizeable fraction of an observatory’s budget. It was therefore very welcome news to read in the Sidereal Times that this had been accomplished that very day! Of course, the production of the Circulars is not without cost, and he noted that Vice President Marsden would report on how this arrangement had in fact been made. West pointed out that, since Commission 6 was outside the new IAU divisional structure, there were special demands on members of Commission 6. It was therefore rather deplorable that a certain fraction of the membership seemed to take only little interest in the commission.

Reporting on the CBAT’s activity during the triennium, Marsden remarked that the downward trend evident in 1994 and 1995 had subsequently been significantly reversed, thanks to the appearance of two spectacular comets and extensive items on γ-ray bursters; furthermore, the 107 supernovae already announced in 1997 already represented a record for a full year. In discussing the financial arrangements whereby it had been possible to place the IAU Circulars freely in the World Wide Web just 160 minutes before the start of the meeting, Marsden noted that, during the preceding year, some 21 percent of the combined income of the CBAT and the Minor Planet Center (MPC) had come from subscribers to the Computer Service. Although there would still be a charge for e-mail delivery of IAU Circulars and use of the Computer Service (CS) on their own computers, the CBAT and the MPC could stand to lose this fraction of their combined income, some 65 percent of which goes for salaries and benefits, principally those of CBAT Associate Director D. W. E. Green and MPC Associate Director G. V. Williams, the latter serving also as webmaster and thus responsible for the Web CS dissemination of the IAU Circulars. What had made the free distribution possible was the anticipation that Williams’ salary for the next four years would instead be paid by a grant from NASA.

A recent editorial in Nature had condemned the use of the IAU Circulars as a forum for the publication of theoretical ideas. Marsden admitted that, although an attempt was made to weed out those sections of the items received that contained theoretical arguments, such publication occasionally happened. Certainly, it seems appropriate to use the Circulars to make a timely prediction that observations thereby inspired might confirm or deny (publication in Nature, for example, being too slow), and some background reasoning may then be necessary, but he agreed that some contributors went too far. It may not be appreciated that many of the items appearing on the IAU Circulars are in fact refereed, but this could always be done more thoroughly. Members of Commission 6 should be, and several of them are an important resource for refereeing activities.

West reported on the Commission 20 vote concerning the uniqueness of comet names. There had in a fact been a tie between those who felt that names should be unique and those for whom it was sufficient that the designations be unique. The status quo would therefore be maintained, at least for the next three years. A second vote had caused the reinstatement of the practice that comets might have three names, rather than the two recommended by the committee set up to produce guidelines for comet naming. S. Nakano remarked on the fact that several amateur astronomers had recently made announcements of their discoveries of alleged comets, novae and supernovae over the internet. It was agreed that this was an unwise development that could deprive the rightful discoverers of credit. If would-be discoverers insist on internet announcement in this way, it was strongly recommended that they also inform the Central Bureau of their finds. It would also be useful if those who do make reports to the Central Bureau were to alert the Bureau to the fact that they had made an internet report. There was a strong feeling that, given the widespread availability of CCD equipment to amateurs, the latter should take some responsibility for making at least some initial confirmation of their claims. Otherwise, the whole concept of a “discovery” could become very blurred. This was somewhat ironic, given the Commission 20 vote on reinstating the practice that comets should be named for up to three discoverers.

Given the very small number of active members of Commission 6, it was agreed that West and Marsden should serve second terms as president and vice president, respectively. K. Aksnes, Nakano and E. Roemer were elected to the Organizing Committee. In the expectation that he would be elected to IAU membership, Green was transferred from consultant to membership status. C. Kouveliotou and M. K. Tsvetkov were also elected as new members of the commission, and G. R. Kastel’ and J. Tichá were appointed as consultants.

The meeting ended with a brief discussion on whether it would be appropriate, in these days of more modern communications, to change the word “telegrams” in the names of both the commission and the Central Bureau. The prevailing view was that there is little wrong with maintaining a harmless tradition. The Bureau in fact still receives telegrams—usually by snailmail—two or three times a year. In any case, some professional contributors frequently refer to the Circulars quite matter-of-factly as “telegrams”, a point that surely—and perhaps appropriately— stresses their urgency.